OpenStudy (anonymous):

HELPPP!!! Find (cos1°+cos2°+…+cos44°)/(sin1°+sin2°+…+sin44°)

OpenStudy (anonymous):

cos1 + cos44 + ...+ cos22 + cos23 ------------------------------ sin1 + sin 44 +...+sin22 + sin23 2cos(45)cos(43)+....+2cos(45)cos(1) ------------------------------ 2sin(45)cos(43)+...+2sin(45)cos1 1

OpenStudy (anonymous):

\( 1+ \sqrt{2} \) ?

OpenStudy (anonymous):

tan 45/2

OpenStudy (anonymous):

not 1, sorry

OpenStudy (anonymous):

as sinc + sind = 2sin(c+d)/2cos(c-d)/2 and cosc + cosd = 2cos(c+d)/2*cos(c-d)/2

OpenStudy (wasiqss):

http://openstudy.com/study#/updates/4f0c9d15e4b084a815fc728d when u done with this do have a look

OpenStudy (anonymous):

tan45 = (2tan45/2)/(1 - tan^2(45/2) 1 -x^2 = 2x note. x = tan45/2

OpenStudy (anonymous):

Ishaan \( \tan \frac {\pi}{8} \) doesn't seem to be the correct answer.

OpenStudy (anonymous):

i did wrong, i know :-(

OpenStudy (anonymous):

cot 45/2 must be it

OpenStudy (anonymous):

x^2 + 2x - 1 + 1 -1 = 0 (x + 1)^2 = 2 x = -1 + sqrt2 x was tan 45/2 henc, cot 45/2 must be 1/(-1+ sqrt2)

OpenStudy (anonymous):

hence*

OpenStudy (anonymous):

actually it;s 2cos(45/2)cos(43/2)+....+2cos(45/2)cos(1/2) ------------------------------ 2sin(45/2)cos(43/2)+...+2sin(45/2)cos1/2 cos(45/2) ( cos43/2 + .. + cos1/2) ------------------------------ sin(45/2)( cos43/2 + ...+cos1/2) which gets down to cot 45/2

OpenStudy (anonymous):

i didn't add (1/2) part in my first post, sorry i am lazy... but now i think you can follow my answer

OpenStudy (anonymous):

kwenisha??

OpenStudy (mertsj):